OpenStudy (anonymous):

The lengths of the sides of a triangle are in the extended ration 3:6:8. The triangles perimeter is 510cm. What are the lengths of the sides?

hartnn (hartnn):

since the sides are in ratio 3:6:8 we can safely assume the sides to be 3x, 6x and 8x now do you know what perimeter is ?

OpenStudy (anonymous):

the length?

hartnn (hartnn):

perimeter means just sum of all sides, so you just add the sides, and equate it to given perimeter of 510 3x+6x +8x =510 can you find 'x' from here ?

OpenStudy (anonymous):

17x?

hartnn (hartnn):

yeah, 17x = 510 so, x=...... ?

OpenStudy (anonymous):

30?

hartnn (hartnn):

yes! thats correct! x= 30 :) then the sides are just 3x = 3*30=... ? 6x = 6*30 =... ? 8x =8*30=.... ?

OpenStudy (anonymous):

90, 180, 240?

hartnn (hartnn):

those are correct side lengths :) any doubts ?

OpenStudy (anonymous):

No. Is that the final answers?

hartnn (hartnn):

yes, thats it! final answer.